OpenStudy (anonymous):

Quick question? Find a ⋅ b. a = 8i + 6j, b = 4i + 5j

OpenStudy (anonymous):

Do the ones with the letters look like this: <32, 30>, or this: 62?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

a.b is a scalar quantity. a.b=(8i+6j)(4i+5j)=(8)(4)+(6)(5)=32+30=62

OpenStudy (anonymous):

Awesome. Thanks. :D

OpenStudy (anonymous):

i.i=i^2=1 i.j=0

OpenStudy (anonymous):

I thought i^2 was -1? Or is that only Algebra?

OpenStudy (anonymous):

i and j are unit vectors. \[i.i=\left| i \right|\left| i \right|\cos 0=(1)(1)(1)=1\]

OpenStudy (anonymous):

that is iota

OpenStudy (anonymous):

iota?

OpenStudy (anonymous):

(iota)^2=-1

OpenStudy (anonymous):

Ohhh. That's the real name of that thing.